A 0.750 mole sample of HBr is introduced into an evacuated (empty) 2.00 L flask and kept at a temperature of 175°C. When the system comes to equilibrium the flask is found to contain 0.206 mole each of H2 and Br2. Determine the value of Kc at 175°C for the equilibrium reaction:

2HBr(g)  ↔ H2(g) + Br2(g)



1
Expert's answer
2021-04-21T03:55:09-0400

80.91 × 0.750= 60.6825 × 2 = 121.365g

0.206 × 2.0157 = 0.415

= 0.206 × 159.808 = 32.92

K = H2(Br)/HBr2

= 0.2062 × 0.206/0.752

= 0.008742/0.5625

= 0.01554

= 1.554 ×10-2